# Deep Learning, Machine Learning - Digital Signal and Image Processing: From Theory to Application

### v1

@@ 1.1 Definition and Applications of Digital Signal Processing
Digital Signal Processing (DSP) refers to the manipulation and analysis of signals after they have been converted to digital form. Signals are a form of data that vary over time, and they can be represented mathematically to convey information.[136] In DSP, we process these signals using algorithms and computational methods in digital computers or specialized processors.
Applications of DSP are vast and span across many fields [155]:
  • Communications: DSP is essential in the modulation and demodulation of signals for transmitting data over long distances, such as in mobile phones, radios, and satellites [155].
  • Audio and Speech Processing: DSP is used in applications like noise cancellation, audio compression, speech recognition, and digital music production [127, 186].
  • Radar and Sonar: DSP helps in the detection and identification of objects by analyzing reflected signals in radar and sonar systems [163].
  • Biomedical Signal Processing: DSP techniques are applied to analyze signals like ECG (Electrocardiogram), EEG (Electroencephalogram), and medical imaging, improving diagnostics and treatment [37].
  • Image Processing: DSP is key in operations such as image enhancement, filtering, compression, and feature extraction in both medical imaging and general photography [58].
@@
  • Precision: Digital systems offer high precision because they use binary numbers, which are less affected by noise and distortion [136, 186].
  • Reusability: Digital systems are programmable, so algorithms can be reused and modified easily without changing hardware [155].
  • Flexibility: DSP allows the implementation of sophisticated algorithms that may be impractical with analog systems.[127]
  • Complexity: Complex algorithms such as adaptive filtering, Fourier transforms, and machine learning techniques can be easily applied in DSP [136].
@@ 1.2 Types of Signals: Discrete and Continuous
A fundamental concept in DSP is the distinction between continuous-time and discrete-time signals [137].
@@ 1.2.1 Continuous-Time Signals
where $t$ represents time, $f$ is the frequency, and the signal is a sine wave [20]. Examples of continuous-time signals include audio signals, temperature measurements, and light intensity variations.
@@
This signal is defined for all $t\in\mathbb{R}$, meaning it exists at every point in time [137]. In practical terms, this could represent an audio tone or a voltage signal.
@@
where $\alpha>0$. This type of signal might represent the charging or discharging of a capacitor in an electrical circuit [136].
@@ 1.2.2 Discrete-Time Signals
where $f$ is the frequency in Hz, and $T_{s}$ is the sampling interval [136].
@@
[155] This could represent temperature measurements at hourly intervals, for example.
@@ 1.3 Python Example: Plotting Signals
In Python, we can easily visualize both continuous and discrete-time signals using libraries like matplotlib and numpy. Here, we will plot a continuous sine wave and its sampled version (discrete signal) [78, 199].
@@ 1.4 From Continuous to Discrete: The Sampling Process
According to the Nyquist-Shannon Sampling Theorem, in order to accurately represent a continuous-time signal in discrete form, the sampling frequency must be at least twice the highest frequency component in the signal [179, 136]:
@@
If the sampling rate is too low (i.e., below the Nyquist rate), a phenomenon called aliasing occurs, where different frequency components become indistinguishable from each other in the sampled data [136].
@@ 1.5 Basic Signal Representation
In this section, we will cover some of the most fundamental signals used in Digital Signal Processing (DSP). These signals form the building blocks for more complex operations and help us understand how systems respond to different inputs. The signals that will be introduced include the unit impulse signal, the unit step signal, sine waves, and complex exponential signals [136, 155].
@@ 1.5.1 Unit Impulse Signal
The unit impulse signal is important because it can be used to represent any discrete-time signal using the principle of superposition. Moreover, it is central to the analysis of systems in the time domain because the response of a system to $\delta[n]$ is known as the impulse response [136].
@@ 1.5.2 Unit Step Signal
The unit step signal is useful for analyzing causal systems, where the output depends only on the current and past inputs, not on future inputs [136].
@@ 1.5.3 Sine and Complex Exponential Signals
where $j$ is the imaginary unit. This form is important because many DSP operations are more easily understood in the frequency domain, where signals are represented as sums of complex exponentials [136].
@@ 1.6 Properties of Linear Time-Invariant (LTI) Systems
LTI systems are foundational in DSP because they simplify analysis and design. Systems with these properties behave in predictable ways, allowing for powerful tools like convolution and the Fourier transform to be used effectively [136, 155].
@@ 1.6.1 Linearity
where $y_{1}[n]$ and $y_{2}[n]$ are the responses of the system to $x_{1}[n]$ and $x_{2}[n]$, respectively. This property makes it easy to analyze systems when the inputs are a combination of simpler signals [136].
@@ 1.6.2 Time Invariance
Time invariance implies that a shift in the input signal results in an identical shift in the output signal. If an LTI system’s response to an input $x[n]$ is $y[n]$, then the response to $x[n-n_{0}]$ (a time-shifted input) will be $y[n-n_{0}]$ (a time-shifted output) [136].
@@
Suppose the system output is delayed by 2 samples relative to the input, i.e., $y[n]=x[n-2]$. This system is time-invariant because if the input is shifted by any amount, the output will be shifted by the same amount.
@@ 1.6.3 Causality and Stability
A system is causal if its output at any time depends only on current and past inputs, not on future inputs. Mathematically, a system is causal if [155]:
@@
A system is stable if it produces bounded outputs for all bounded inputs. In other words, if the input signal is bounded, the output signal must also be bounded. This can be written as [136]:
@@
The system $y[n]=0.5y[n-1]+x[n]$ is causal because the output at time $n$ depends only on the current and previous inputs. It is also stable as long as the feedback coefficient (0.5 in this case) is less than 1, ensuring that the output does not grow unbounded.
@@ 1.7.1 Discrete-Time Convolution
Here, $x[k]$ is the input signal, $h[n-k]$ is the system’s response, and the sum is computed over all values of $k$. The result $y[n]$ is a new signal that represents the combined effect of $x[n]$ passing through the system with response $h[n]$.
@@ 1.7.3 Correlation
Correlation is another key concept in signal processing, used to measure the similarity between two signals. It is closely related to convolution, but instead of sliding one signal over the other as in convolution, correlation compares the signals as they are [136, 155].
@@ 1.8.1 Nyquist Sampling Rate
The Nyquist-Shannon Sampling Theorem is a foundational concept in digital signal processing. It states that a continuous signal can be perfectly reconstructed from its discrete samples if the sampling rate is at least twice the highest frequency component present in the signal. This critical frequency is known as the Nyquist rate, and sampling below this rate can lead to a phenomenon known as aliasing [136, 179].
@@
[…]
@@ 1.8.2 Signal Reconstruction and Anti-Aliasing Filters
Once a signal is sampled, it can be reconstructed by passing the samples through a low-pass filter. This filter removes the high-frequency components that could cause aliasing. Before sampling, an anti-aliasing filter is applied to ensure that no frequency components above half the sampling rate are present in the signal [136].
@@ 2.1 Introduction to Fourier Transform
The Fourier Transform is a fundamental mathematical tool used in signal processing to decompose a signal into its constituent sinusoidal components, each with a different frequency, amplitude, and phase [136, 20]. In the case of continuous signals, we use the Continuous Fourier Transform (CFT). However, in the digital world, we deal with discrete-time signals, and thus we need to use the Discrete-Time Fourier Transform (DTFT) to analyze the frequency content of such signals.
@@ 2.2 Definition of DTFT
The DTFT transforms a discrete-time signal $x[n]$, which is defined only at specific time instances, into a continuous function of frequency $X(e^{j\omega})$. This gives us insight into how the energy of the signal is distributed across different frequencies [136].
@@ 2.2.1 Key Concepts of DTFT
  • Complex Exponentials: In the expression $e^{-j\omega n}$, we see the use of a complex exponential. This is a key aspect of Fourier Transforms, as they represent sinusoidal functions (sines and cosines) in complex form [136].
  • Frequency $\omega$: The variable $\omega$ represents the angular frequency, and it is continuous. For practical purposes, we often normalize it to range between $-\pi$ and $\pi$ radians.
  • Summation Over All Time Indices: The summation is over all possible time indices $n$, from $\infty$- ∞ to $\infty$. This captures the entire signal’s behavior in both the past and future [136].
@@
Therefore, the DTFT of the unit impulse is constant across all frequencies [136].
@@ 2.3 Properties of DTFT
The DTFT has several important properties that make it useful in signal processing. Understanding these properties allows us to work with complex systems more effectively[136, 155].
@@ 2.3.1 Linearity
where $a$ and $b$ are constants. This property is very useful when combining multiple signals or designing systems that process multiple inputs [119].
@@ 2.3.2 Time Shifting
This shows that time shifts correspond to phase shifts in the frequency domain [137].
@@ 2.3.3 Frequency Shifting
This property is used in modulation techniques, where signals are shifted to higher or lower frequencies [20].
@@ 2.3.4 Convolution Theorem
where * denotes convolution. This is extremely useful in filtering applications [127, 58].
@@ 2.6 Properties of Discrete-Time Fourier Transform (DTFT)
The Discrete-Time Fourier Transform (DTFT) is a fundamental tool in Digital Signal Processing (DSP) that allows us to analyze signals in the frequency domain. The DTFT is particularly useful for periodic signals and systems analysis, providing insight into how signals behave across different frequencies. In this section, we will explore the key properties of the DTFT, such as linearity, time-shifting, and the convolution theorem, and provide detailed examples to make these concepts clear [136, 155, 127].
@@ 2.6.1 Linearity
where $X_{1}(e^{j\omega})$ and $X_{2}(e^{j\omega})$ are the DTFTs of $x_{1}[n]$ and $x_{2}[n]$, respectively [136, 20].
@@ 2.6.2 Time-Shifting Property
The phase shift introduced in the frequency domain is proportional to the time shift in the time domain [136, 119].
@@ 2.6.3 Convolution Theorem
This property is especially useful in filtering and signal processing, where convolution is a common operation, and working in the frequency domain can simplify the computation [127, 58].
@@ 2.7 Inverse Discrete-Time Fourier Transform (Inverse DTFT)
[…]
In this equation, $X(e^{j\omega})$ is the frequency domain representation (the DTFT) of the signal $x[n]$, and the term $e^{j\omega n}$ is the inverse Fourier kernel that allows us to reconstruct the original time-domain signal. The integral sums up the frequency components of the signal from $-\pi$ to $\pi$, which corresponds to the full frequency range for a discrete-time signal [137].
@@ 2.8 DTFT for Periodic Signals
For periodic discrete-time signals, the DTFT behaves differently than for non-periodic signals. A periodic signal in the time domain results in a discrete set of impulses in the frequency domain, reflecting the fact that the signal consists of harmonics or multiples of a fundamental frequency [136]. This is due to the periodic nature of the signal, which leads to a periodic repetition of the signal’s frequency components.
@@
Here, $X[k]$ are the Fourier series coefficients, and the frequency domain consists of impulses at intervals of $2\pi/N$ [155, 137].
@@ 3.1 Definition of Discrete Fourier Transform
The Discrete Fourier Transform (DFT) is a fundamental mathematical tool in digital signal processing. It is a discrete version of the Fourier transform that applies to finite-length sequences. Unlike the DTFT, which is applied to infinite sequences and provides a continuous frequency spectrum, the DFT is used for signals of finite length and yields a discrete set of frequency components [136, 155, 20, 127, 156].
@@
  • $N$ is the length of the input sequence,
@@
The DFT is a powerful tool because it converts a finite sequence of time-domain data into a finite sequence of frequency-domain data. This transformation allows us to analyze the frequency content of the signal, apply filters, and perform various signal processing operations [127, 136, 156].

### v2

@@
   Xinyuan Song
Emory University
xsong30@emory.edu
@@ 1.1 Definition and Applications of Digital Signal Processing
Digital Signal Processing (DSP) refers to the manipulation and analysis of signals after they have been converted to digital form. Signals are a form of data that vary over time, and they can be represented mathematically to convey information.[oppenheim2010dsp] In DSP, we process these signals using algorithms and computational methods in digital computers or specialized processors.
Applications of DSP are vast and span across many fields [proakis1996digital]:
  • Communications: DSP is essential in the modulation and demodulation of signals for transmitting data over long distances, such as in mobile phones, radios, and satellites [proakis1996digital].
  • Audio and Speech Processing: DSP is used in applications like noise cancellation, audio compression, speech recognition, and digital music production [mitra2002digital, smith1997scientist].
  • Radar and Sonar: DSP helps in the detection and identification of objects by analyzing reflected signals in radar and sonar systems [richards2014fundamentals].
  • Biomedical Signal Processing: DSP techniques are applied to analyze signals like ECG (Electrocardiogram), EEG (Electroencephalogram), and medical imaging, improving diagnostics and treatment [cohen2019biomedical].
  • Image Processing: DSP is key in operations such as image enhancement, filtering, compression, and feature extraction in both medical imaging and general photography [gonzalez2002digital].
@@
  • Precision: Digital systems offer high precision because they use binary numbers, which are less affected by noise and distortion [oppenheim2010dsp, smith1997scientist].
  • Reusability: Digital systems are programmable, so algorithms can be reused and modified easily without changing hardware [proakis1996digital].
  • Flexibility: DSP allows the implementation of sophisticated algorithms that may be impractical with analog systems.[mitra2002digital]
  • Complexity: Complex algorithms such as adaptive filtering, Fourier transforms, and machine learning techniques can be easily applied in DSP [oppenheim2010dsp].
@@ 1.2 Types of Signals: Discrete and Continuous
A fundamental concept in DSP is the distinction between continuous-time and discrete-time signals [oppenheim1997signals].
@@ 1.2.1 Continuous-Time Signals
where $t$ represents time, $f$ is the frequency, and the signal is a sine wave [bracewell2000fourier]. Examples of continuous-time signals include audio signals, temperature measurements, and light intensity variations.
@@
This signal is defined for all $t\in\mathbb{R}$, meaning it exists at every point in time [oppenheim1997signals]. In practical terms, this could represent an audio tone or a voltage signal.
@@
where $\alpha>0$. This type of signal might represent the charging or discharging of a capacitor in an electrical circuit [oppenheim2010dsp].
@@ 1.2.2 Discrete-Time Signals
where $f$ is the frequency in Hz, and $T_{s}$ is the sampling interval [oppenheim2010dsp].
@@
[proakis1996digital] This could represent temperature measurements at hourly intervals, for example.
@@ 1.3 Python Example: Plotting Signals
In Python, we can easily visualize both continuous and discrete-time signals using libraries like matplotlib and numpy. Here, we will plot a continuous sine wave and its sampled version (discrete signal) [hunter2007matplotlib, van2011numpy].
@@ 1.4 From Continuous to Discrete: The Sampling Process
According to the Nyquist-Shannon Sampling Theorem, in order to accurately represent a continuous-time signal in discrete form, the sampling frequency must be at least twice the highest frequency component in the signal [shannon1949communication, oppenheim2010dsp]:
@@
If the sampling rate is too low (i.e., below the Nyquist rate), a phenomenon called aliasing occurs, where different frequency components become indistinguishable from each other in the sampled data [oppenheim2010dsp].
@@ 1.5 Basic Signal Representation
In this section, we will cover some of the most fundamental signals used in Digital Signal Processing (DSP). These signals form the building blocks for more complex operations and help us understand how systems respond to different inputs. The signals that will be introduced include the unit impulse signal, the unit step signal, sine waves, and complex exponential signals [oppenheim2010dsp, proakis1996digital].
@@ 1.5.1 Unit Impulse Signal
The unit impulse signal is important because it can be used to represent any discrete-time signal using the principle of superposition. Moreover, it is central to the analysis of systems in the time domain because the response of a system to $\delta[n]$ is known as the impulse response [oppenheim2010dsp].
@@ 1.5.2 Unit Step Signal
The unit step signal is useful for analyzing causal systems, where the output depends only on the current and past inputs, not on future inputs [oppenheim2010dsp].
@@ 1.5.3 Sine and Complex Exponential Signals
where $j$ is the imaginary unit. This form is important because many DSP operations are more easily understood in the frequency domain, where signals are represented as sums of complex exponentials [oppenheim2010dsp].
@@ 1.6 Properties of Linear Time-Invariant (LTI) Systems
LTI systems are foundational in DSP because they simplify analysis and design. Systems with these properties behave in predictable ways, allowing for powerful tools like convolution and the Fourier transform to be used effectively [oppenheim2010dsp, proakis1996digital].
@@ 1.6.1 Linearity
where $y_{1}[n]$ and $y_{2}[n]$ are the responses of the system to $x_{1}[n]$ and $x_{2}[n]$, respectively. This property makes it easy to analyze systems when the inputs are a combination of simpler signals [oppenheim2010dsp].
@@ 1.6.2 Time Invariance
Time invariance implies that a shift in the input signal results in an identical shift in the output signal. If an LTI system’s response to an input $x[n]$ is $y[n]$, then the response to $x[n-n_{0}]$ (a time-shifted input) will be $y[n-n_{0}]$ (a time-shifted output) [oppenheim2010dsp].
@@
Suppose the system output is delayed by 2 samples relative to the input, i.e., y[n]=x[n2]y[n]=x[n-2]. This system is time-invariant because if the input is shifted by any amount, the output will be shifted by the same amount.
@@ 1.6.3 Causality and Stability
A system is causal if its output at any time depends only on current and past inputs, not on future inputs. Mathematically, a system is causal if [proakis1996digital]:
@@
A system is stable if it produces bounded outputs for all bounded inputs. In other words, if the input signal is bounded, the output signal must also be bounded. This can be written as [oppenheim2010dsp]:
@@
The system y[n]=0.5y[n1]+x[n]y[n]=0.5y[n-1]+x[n] is causal because the output at time $n$ depends only on the current and previous inputs. It is also stable as long as the feedback coefficient (0.5 in this case) is less than 1, ensuring that the output does not grow unbounded.
@@ 1.7.1 Discrete-Time Convolution
Here, $x[k]$ is the input signal, h[nk]h[n-k] is the system’s response, and the sum is computed over all values of $k$. The result $y[n]$ is a new signal that represents the combined effect of $x[n]$ passing through the system with response $h[n]$.
@@ 1.7.3 Correlation
Correlation is another key concept in signal processing, used to measure the similarity between two signals. It is closely related to convolution, but instead of sliding one signal over the other as in convolution, correlation compares the signals as they are [oppenheim2010dsp, proakis1996digital].
@@ 1.8.1 Nyquist Sampling Rate
The Nyquist-Shannon Sampling Theorem is a foundational concept in digital signal processing. It states that a continuous signal can be perfectly reconstructed from its discrete samples if the sampling rate is at least twice the highest frequency component present in the signal. This critical frequency is known as the Nyquist rate, and sampling below this rate can lead to a phenomenon known as aliasing [oppenheim2010dsp, shannon1949communication].
@@
$\text{Sampling rate}\geq 2\times 500\text{ Hz}=1000\text{ Hz}$
@@ 1.8.2 Signal Reconstruction and Anti-Aliasing Filters
Once a signal is sampled, it can be reconstructed by passing the samples through a low-pass filter. This filter removes the high-frequency components that could cause aliasing. Before sampling, an anti-aliasing filter is applied to ensure that no frequency components above half the sampling rate are present in the signal [oppenheim2010dsp].
@@ 2.1 Introduction to Fourier Transform
The Fourier Transform is a fundamental mathematical tool used in signal processing to decompose a signal into its constituent sinusoidal components, each with a different frequency, amplitude, and phase [oppenheim2010dsp, bracewell2000fourier]. In the case of continuous signals, we use the Continuous Fourier Transform (CFT). However, in the digital world, we deal with discrete-time signals, and thus we need to use the Discrete-Time Fourier Transform (DTFT) to analyze the frequency content of such signals.
@@ 2.2 Definition of DTFT
The DTFT transforms a discrete-time signal $x[n]$, which is defined only at specific time instances, into a continuous function of frequency $X(e^{j\omega})$. This gives us insight into how the energy of the signal is distributed across different frequencies [oppenheim2010dsp].
@@ 2.2.1 Key Concepts of DTFT
  • Complex Exponentials: In the expression $e^{-j\omega n}$, we see the use of a complex exponential. This is a key aspect of Fourier Transforms, as they represent sinusoidal functions (sines and cosines) in complex form [oppenheim2010dsp].
  • Frequency $\omega$: The variable $\omega$ represents the angular frequency, and it is continuous. For practical purposes, we often normalize it to range between $\pi$ and $\pi$ radians.
  • Summation Over All Time Indices: The summation is over all possible time indices $n$, from -$\infty$ to $\infty$. This captures the entire signal’s behavior in both the past and future [oppenheim2010dsp].
@@
Therefore, the DTFT of the unit impulse is constant across all frequencies [oppenheim2010dsp].
@@ 2.3 Properties of DTFT
The DTFT has several important properties that make it useful in signal processing. Understanding these properties allows us to work with complex systems more effectively[oppenheim2010dsp, proakis1996digital].
@@ 2.3.1 Linearity
where $a$ and $b$ are constants. This property is very useful when combining multiple signals or designing systems that process multiple inputs [lyons2010understanding].
@@ 2.3.2 Time Shifting
This shows that time shifts correspond to phase shifts in the frequency domain [oppenheim1997signals].
@@ 2.3.3 Frequency Shifting
This property is used in modulation techniques, where signals are shifted to higher or lower frequencies [bracewell2000fourier].
@@ 2.3.4 Convolution Theorem
where * denotes convolution. This is extremely useful in filtering applications [mitra2002digital, gonzalez2002digital].
@@ 2.6 Properties of Discrete-Time Fourier Transform (DTFT)
The Discrete-Time Fourier Transform (DTFT) is a fundamental tool in Digital Signal Processing (DSP) that allows us to analyze signals in the frequency domain. The DTFT is particularly useful for periodic signals and systems analysis, providing insight into how signals behave across different frequencies. In this section, we will explore the key properties of the DTFT, such as linearity, time-shifting, and the convolution theorem, and provide detailed examples to make these concepts clear [oppenheim2010dsp, proakis1996digital, mitra2002digital].
@@ 2.6.1 Linearity
where $X_{1}(e^{j\omega})$ and $X_{2}(e^{j\omega})$ are the DTFTs of $x_{1}[n]$ and $x_{2}[n]$, respectively [oppenheim2010dsp, bracewell2000fourier].
@@ 2.6.2 Time-Shifting Property
The phase shift introduced in the frequency domain is proportional to the time shift in the time domain [oppenheim2010dsp, lyons2010understanding].
@@ 2.6.3 Convolution Theorem
This property is especially useful in filtering and signal processing, where convolution is a common operation, and working in the frequency domain can simplify the computation [mitra2002digital, gonzalez2002digital].
@@ 2.7 Inverse Discrete-Time Fourier Transform (Inverse DTFT)
$x[n]=\frac{1}{2\pi}\int_{-\pi}^{\pi}X(e^{j\omega})e^{j\omega n}d\omega$
In this equation, $X(e^{j\omega})$ is the frequency domain representation (the DTFT) of the signal $x[n]$, and the term $e^{j\omega n}$ is the inverse Fourier kernel that allows us to reconstruct the original time-domain signal. The integral sums up the frequency components of the signal from $\pi$ to $\pi$, which corresponds to the full frequency range for a discrete-time signal [oppenheim1997signals].
@@ 2.8 DTFT for Periodic Signals
For periodic discrete-time signals, the DTFT behaves differently than for non-periodic signals. A periodic signal in the time domain results in a discrete set of impulses in the frequency domain, reflecting the fact that the signal consists of harmonics or multiples of a fundamental frequency [oppenheim2010dsp]. This is due to the periodic nature of the signal, which leads to a periodic repetition of the signal’s frequency components.
@@
Here, $X[k]$ are the Fourier series coefficients, and the frequency domain consists of impulses at intervals of $2\pi/N$ [proakis1996digital, oppenheim1997signals].
@@ 3.1 Definition of Discrete Fourier Transform
The Discrete Fourier Transform (DFT) is a fundamental mathematical tool in digital signal processing. It is a discrete version of the Fourier transform that applies to finite-length sequences. Unlike the DTFT, which is applied to infinite sequences and provides a continuous frequency spectrum, the DFT is used for signals of finite length and yields a discrete set of frequency components [oppenheim2010dsp, proakis1996digital, bracewell2000fourier, mitra2002digital, rader1968discrete].
@@
  • NN is the length of the input sequence,
@@
The DFT is a powerful tool because it converts a finite sequence of time-domain data into a finite sequence of frequency-domain data. This transformation allows us to analyze the frequency content of the signal, apply filters, and perform various signal processing operations [mitra2002digital, oppenheim2010dsp, rader1968discrete].